# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v2

@@ 3.4 Ablation Study: Removing High Quality Small Dataset
To test whether our empirical gains are conditioned on the traces generated by our procedure and on a highly optimized small dataset (muennighoff2025s1_simpletesttime), we fine-tune on a public dataset that already provides sufficient reasoning traces per question: the 93k math set of openr1 (2-4 traces per question). Because this dataset has been successful for fine-tuning Qwen2.5-Math-7B, we adopt that base model and compare SSFT against SFT trained on all available traces. Details on the hyperparameters are in Appendix A.3.4. Addressing RQ5, Table 2 shows consistent improvements in both $\mathrm{Pass@}1$ and $\mathrm{Cons@}32$. The results indicate the SSFT is effective for larger sized public dataset with less diverse reasoning traces.
@@
Table 2: Performance of SSFT versus SFT trained solely on publicly available distillation targets. The setup uses the 93k math questions from openr1 with Qwen2.5-Math-7B as the base model. SFT-OpenR1-93k-7B uses the same distillation targets as SFT-mixed-distill-32B in Table 1.
@@ 4 Related work
Test-time Scaling. There has been a surge of work fine-tuning LLMs to reason longer, using reinforcement learning for frontier models (o1; shao2024deepseekmathpushinglimitsmathematical; xAI2025; yang2025qwen3) and supervised fine-tuning for smaller ones (muennighoff2025s1_simpletesttime; hu2025distillationwhydistillation). These methods enable LLMs to improve reasoning by allocating more test-time compute to sequential, iterative refinement such as self-reflection (guo2025deepseekR1; liu2025understanding_r1zero). However, extended sequential reasoning can be more sensitive to the order of reasoning steps and may result in failures (chen2024premiseOrderMatters), and performance can start to degrade beyond a certain length due to “overthinking” (ghosal2025doesoverthinkinghelp). Our goal is to study the effective use of diverse reasoning traces to fine-tune small language models, essential for agentic AI (belcak2025small). Related to our work on aggregating diverse reasoning modes, yu2025chainofreasoning_towardsunifedmathematical; zheng2025learningtoreasonviaMixtureOfThought proposed incorporating three specific reasoning modes: natural language, code-based, and symbolic reasoning.
@@
Parallel Reasoning. Parallel scaling methods such as self-consistency  (wang2022selfconsistency) and Best-of-N (best_of_n) improve LLM performance by generating multiple reasoning paths in parallel and aggregating them. These methods fundamentally require choosing a temperature that can generate diverse reasoning paths, but a recent theoretical work shows that increasing temperature can sometimes fail to increase diversity if language models are not trained towards coverage (verine2025improvingwhentemperaturefailschangetheloss). Other search-based methods such as Monte Carlo tree search (MCTS) (zhang2024accessingmontecarlo) and Tree of Thoughts (ToT) (yao2023treeofthoughts) apply heuristic-guided search with an external verifier to do more deliberate search to increase the coverage. However, their dependence on heuristics and domain-specific knowledge can limit their applicable tasks. Regarding training LLMs with parallel reasoning traces,  yang2025multiverse proposes training with parallel CoTs decomposed from sequential CoTs, and our concurrent work wen2025parathinker proposes to train with multiple reasoning traces distilled from teacher models. These works show native parallel scaling can surpass sequential scaling within certain token limits. Although we also show improvements in reasoning performance by simply fine-tuning with parallel reasoning traces, we empirically highlighted the limitation of manually assigning parallel think tags to target reasoning traces. If human assignment or random enforcement routes mismatched traces to a tag with a unique reasoning mode learned, the tag’s semantics blur due to bias in the ordering of parallel reasoning traces. This prevents tags from initiating distinct reasoning modes. We are the only work that leverages parallel think tags to construct a set-prediction loss via bipartite-matching optimization between tags and reasoning traces, turning the tags into global forking tokens. Empirically, we also improve $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ over baselines fine-tuned with parallel reasoning traces on the same dataset, whether on subsets or the full set. Another recent work (zheng2025parallelr1towardsparallelthinkingviaRL) encourages the use of parallel thinking tags by slightly upweighting rewards for correct trajectories that employ such tags during reinforcement learning. However, the cold-start SFT that learns parallel thinking from diverse reasoning traces does not incorporate a set-prediction loss, which is necessary to prevent collapse in reasoning diversity.
@@
Set-based Global Loss in Deep Learning. DETR introduces end-to-end object detection with a set global loss (carion2020end_detr; owl_vit), whose success in parallel bounding-box prediction inspires our approach. We are the first to extend this to language modeling: While DETR predicts a set of single tokens in parallel to match a list of bounding boxes, we predict a set of sequential thinking tokens, initiated by global forking tokens, and evaluate the matchings using autoregressive losses. We adapt set-based loss with autoregressive (AR) models, rather than diffusion-based models (zhao2025d1; nie2025largediffusionmodels; DDPM), because AR models achieve superior reasoning performance. Several innovations are needed to make this loss practical for parallel reasoning with AR models. For example, to populate the cost matrix we need $\mathrm{N}\times\mathrm{M}$ forward passes, one for each combination of $\mathrm{N}$ think tags and $\mathrm{M}$ reasoning traces, whereas DETR requires only $\mathrm{N}$ forward passes for its $\mathrm{N}$ queries, since the same bounding-box predictions can be reused to compute matching costs against different ground-truth boxes. We therefore emphasize efficient training implementation in our extension, increasing training time by only $6.6\%$ compared to standard SFT (Appendix A.1).
@@ 5 Conclusion
In this work, we demonstrate that diverse reasoning traces can be leveraged to learn global forking tokens which serve as prompts to initiate distinct reasoning modes that are both diverse and accurate. We propose Set Supervised Fine-Tuning (SSFT), which employs bipartite matching between reserved global forking tokens and diverse reasoning traces to compute a set-based language modeling loss. We show that models trained with SSFT yield improvements in both $\mathrm{Pass@}k$ and $\mathrm{Cons@}k$ accuracy over various $k$ compared to standard SFT with temperature scaling. The proposed training method also improves the usage of multi-target data and avoids collapsing distinct reasoning modes, which is especially helpful in the context of distilling from multiple teacher models. In future work, we plan to further investigate how performance scales with a larger number of distillation targets and reserved forking tokens, which jointly determine the size of the underlying bipartite graph. In addition, we will apply RFT to a model pre-trained with SSFT to incentivize optimal use of the learned global forking tokens across a broader suite of tasks without requiring additional high-quality reasoning traces.
Acknowledgment
This research was supported by compute resources provided by Amazon Web Services.
@@ Appendix A Appendix
In this appendix, we provide details omitted from the main text.
A.1 Algorithm: SSFT Implementation

### v3

@@ 4 Robustness and Ablation Studies
Robustness to Fine-Tuning Data Quality. To test whether our gains depend on the traces generated by our procedure and on a small high-quality dataset (Muennighoff et al., 2025), we fine-tune on a public dataset that already provides 2-4 reasoning traces per question: the 93k math set of Face (2025). Because this dataset has been successful for fine-tuning Qwen2.5-Math-7B, we adopt that base model and compare SSFT against SFT trained on all available traces. Details are in Appendix A.9.4. Addressing RQ5, Table 3 shows consistent improvements in both $\mathrm{Pass@}1$ and $\mathrm{Cons@}32$. The results indicate SSFT is effective for larger public dataset with less diverse traces.
Table 3: Performance of SSFT versus SFT trained solely on publicly available distillation targets. The setup uses the 93k math questions from Face (2025) with Qwen2.5-Math-7B as the base model. SFT-OpenR1-93k-7B uses the same distillation targets as SFT-mixed-distill-32B in Table 1.
@@
Robustness to Model Family & Size. In Appendix A.5, SSFT also shows gains with Qwen3-4B-Base and Llama3.1-8B-Instruct, though the improvements are larger for the bigger Qwen models.
Impact of Matching Length. We choose $\mathrm{L}=\frac{\text{max\_seq\_len}}{\mathrm{M}\mathrm{N}}\approx 1000$ because it is roughly the largest value that allows $\mathrm{N}\mathrm{M}$ matching costs to be computed in a single forward pass without extra VRAM. Figure 7 shows this adds very minimal computation because backpropagation and storing activations still dominate the training time. Note that in our setting, $\mathrm{L}=1000$ was sufficient for maximally differentiating complex reasoning modes and preserving high-accuracy traces, as reflected by $\mathrm{Pass@}1$ and diversity in $\mathrm{Cons@}k$ in Figure 7. Increasing $\mathrm{L}$ does not hurt $\mathrm{Pass@}1$ or $\mathrm{Cons@}32$. When additional compute time permits, $\mathrm{L}$ can be increased further and stopped once performance plateaus.
Figure 7: Left: Impact on training time. Middle: Impact on $\mathrm{Pass@}1$. Right: Impact on $\mathrm{Cons@}32$. Ablation on the number of tokens $\mathrm{L}$. This is a hyperparameter when computing the matching cost in Eq. 2. Setting $\mathrm{L}=0$ reduces to SSFT with random matching. For each $\mathrm{L}$, we apply SSFT with matching $(\mathrm{N}=6,\mathrm{M}=4)$ to fine-tune Qwen2.5-32B-Instruct on s1k-4mixed-reasoning and evaluate training time, Pass@1 on two challenging reasoning tasks, and $\mathrm{Cons@}32$ on the same tasks.
@@ 5 Related work
Test-time Scaling. There has been a surge of work fine-tuning LLMs to reason longer, using reinforcement learning for frontier models (OpenAI, 2024; Shao et al., 2024; xAI, 2025; Yang et al., 2025a) and supervised fine-tuning for smaller ones (Muennighoff et al., 2025; Hu et al., 2025). These methods enable LLMs to improve reasoning by allocating more test-time compute to sequential, iterative refinement such as self-reflection (Guo et al., 2025; Liu et al., 2025a). However, extended sequential reasoning can be more sensitive to the order of reasoning steps and may result in failures (Chen et al., 2024b), and performance can start to degrade beyond a certain length due to “overthinking” (Ghosal et al., 2025). Our goal is to study the effective use of diverse reasoning traces to fine-tune small language models, essential for agentic AI (Belcak et al., 2025).
@@
Parallel Reasoning. Parallel scaling methods such as self-consistency  (Wang et al., 2022) and Best-of-N (Lightman et al., 2023) improve LLM performance by generating multiple reasoning paths in parallel and aggregating them. These methods fundamentally require choosing a temperature that can generate diverse reasoning paths, but a recent theoretical work shows that increasing temperature can sometimes fail to increase diversity if language models are not trained towards coverage (Verine et al., 2025). Other search-based methods such as Monte Carlo tree search (MCTS) (Zhang et al., 2024) and Tree of Thoughts (ToT) (Yao et al., 2023) apply heuristic-guided search with an external verifier to do more deliberate search to increase the coverage (Yao et al., 2023; Zhang et al., 2024). However, their dependence on heuristics and domain-specific knowledge can limit their applicable tasks. Regarding training LLMs with parallel reasoning traces,  Yang et al. (2025b) proposes training with parallel CoTs decomposed from sequential CoTs, and our concurrent work Wen et al. (2025) proposes to train with multiple reasoning traces distilled from teacher models. These works show native parallel scaling can surpass sequential scaling within certain token limits. However, we aim to show that training on diverse distilled traces with our set language modeling loss enables the model to learn global forking tokens that trigger distinct reasoning modes, improving $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ over baselines fine-tuned on the same dataset, whether on subsets or the full set.
@@
Set-based Global Loss in Deep Learning. DETR introduces end-to-end object detection with a set global loss (Carion et al., 2020; Minderer et al., 2022), whose success in parallel bounding-box prediction inspires our approach. We are the first to extend this to language modeling: while DETR predicts a set of tokens in parallel to match a list of bounding boxes, we predict a set of sequential reasoning paths initiated by global forking tokens and assess the matchings based on autoregressive losses. We adapt the set-based loss with autoregressive (AR) models, rather than diffusion-based models (Nie et al., 2025), because AR models achieve superior reasoning performance.
@@ 6 Conclusion
We show that diverse reasoning traces can be leveraged to learn global forking tokens that globally steer diverse and accurate reasoning modes. We propose Set Supervised Fine-Tuning (SSFT), which employs bipartite matching between such control tokens and traces to compute a set language modeling loss. We visualize that SSFT preserves distinct reasoning modes that SFT collapses, and produces emergent global forking tokens that Global Forking Policy Optimization leverages to incentivize complex reasoning. The resulting models consistently outperform their SFT counterparts with GRPO, improving $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ across math reasoning benchmarks and $\mathrm{Pass@}1$ on execution-based code generation benchmarks. For future work, we plan to scale up the underlying bipartite graph and explore more general settings beyond multi-teacher distillation.
Acknowledgements
The authors thank Michael Zhang, Silviu Pitis, Andrew Li, and Elena Zhang at the University of Toronto for their valuable feedback. We also acknowledge Amazon Web Services for providing the computing resources that made this research possible.
@@ Appendix A Appendix
A.1 LLM Usage in Paper
We used ChatGPT (OpenAI, GPT-5) in September 2025 for occasional language polishing only. This is done when we really wanted to make sure there are no grammatical errors in a few sentences. No text, code, experiment results, or figures were generated by the LLM. We made our own hypothesis, completed all technical content on our own, and made our conclusions. We also verified all outputs by ourselves. The occasional grammar check is done by typing into LLM chatbox.
A.2 Learned Matchings between Global Forking Tokens and Traces
A.2.1 How to Choose the Global Forking Token for Pass@1
A.2.2 More Visualizations on Learned Matchings
(a) Visualization of learned matchings after SSFT with random matching. The fully connected graph shows no correlations were learned
(b) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
(c) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
Figure 8: These learned matching visualizations are obtained by connecting edges in the subset of configurations $\{\bm{\sigma}_{k}\}_{k=1}^{p}$ that still accumulate mass towards the end of training in Figure 6. These models are fine-tuned using the same GPT-OSS-high, GPT-OSS-medium, R1, and Gemini traces for each question, so we can better interpret the learned matchings
A.3 Global Forking Policy Optimization for $\mathrm{Pass@}1$ Inference by Sampling ${\textnormal{g}}^{(i)}$.
For single-path generation, we prompt the ${\textnormal{g}}^{(i)}$ most likely to correlate with a complex reasoning mode using a graph-based heuristic that selects the ${\textnormal{g}}^{(i)}$ matched to the largest number of distinct reasoning traces. Since theoretically justifying any single heuristic is difficult, we apply a small amount of RL to learn the optimal distribution over global forking tokens conditioned on each question. Using cold-start SFT before RL is common practice in LLM post-training (Guo et al., 2025), and recent work shows that RLVR alone does not incentivize new capabilities (Yue et al., 2025; Liu et al., 2025b).
Setup. We use 1,280 questions from DAPO-Math-17k and run GRPO for 10 steps using only the policy gradients of ${\textnormal{g}}^{(i)}$. These are referred to as GFPO. We also let a baseline optimize full tokens even though that is much more computationally expensive. The base SFT models include SSFT-32B, SSFT-32B (random $\bm{\sigma}$), SFT-mixed-distill-32B-tags, trained using s1k-4mixed-reasoning. Hyperparameters are in Table 9.
Results. In Table 4, we show that SSFT instills diverse, high-accuracy reasoning modes without mode collapse, and that merely 10 steps of GFPO can shape the output distribution of ${\textnormal{g}}^{(i)}$ conditioned on $\mathbf{x}$}. By comparing against baselines with RL, we see that the gains come from the synergy between GFPO and the global forking tokens learned by SSFT with a principled set loss, rather than from spurious rewards (Shao et al., 2025). In our setting, SFT-GRPO that optimizes all tokens still cannot match SSFT-GFPO, despite using substantially more compute.
Table 4: Comparison of SSFT and baselines on three math reasoning tasks after further RL fine-tuning only on global forking tokens, denoted as GFPO. We also include an expensive GRPO baseline that optimizes the full sequences. SFT/SSFT uses s4k-mixed-reasoning data, and GRPO/GFPO uses only 1280 questions from DAPO-Math-17K (Yu et al., 2025). The optimal <think i>  for each question is sampled by the models in this evaluation. By avoiding reasoning mode collapse observed in standard SFT-mixed-distill and SSFT (random $\bm{\sigma}$) with random matching, SSFT preserves complex, high-accuracy reasoning modes. Standard GRPO that applies gradients to full sequences does not improve performance and can even lead to a slight drop in this small-scale RL setting.
AIME 2024 AIME 2025 MATH-500 $\text{LCB(v5)}^{*}$
Pass@1: Average performance of 32 individual generations (AIME24/25, MATH-500)
SFT-mixed-distill-32B-tags-GRPO 58.85 52.40 88.85 37.13
SFT-mixed-distill-32B-tags-GFPO 59.80 54.06 88.98 37.72
SSFT-32B (random $\bm{\sigma}$)- GFPO 59.58 53.96 89.87 36.53
SSFT-32B- GFPO 64.22 58.80 89.90 42.10
A.4 Robustness to Fine-tuning with Code Generation Traces
We investigate the advantage of SSFT over SFT when the fine-tuning dataset consists of traces from open-ended domains, such as code generation with unit test cases. We study whether SSFT can improve $\mathrm{Pass@}1$ on competitive coding by instilling specialized coding reasoning modes without mode collapse.
Setup. We select 1K coding questions from Open-Thoughts (Guha et al., 2025), use DeepSeek-R1 to generate four reasoning traces per question to construct code1k-4mixed-reasoning. We apply SSFT on this dataset, and then GFPO on 1k questions from Intellect-2-RL (Team et al., 2025). At test time, each model samples <think i>  for each question. For evaluation, we use LiveCodeBench(v5) and assess whether the models can generate valid code that passes unit tests when executed. Although the evaluation still relies on a verifiable metric that avoids evaluation bias, the resulting code is far more open-ended than in math tasks, where only a final numeric answer is extracted and scored. We still included the math tasks for OOD evaluation.
Results. Table 5 reports the full comparison between SSFT and baseline methods under fine-tuning with code-generation traces. SSFT outperforms the baselines on LiveCodeBench and shows stronger generalization to the math domain, despite being fine-tuned only on code-generation traces. These results show the robustness of SSFT to using traces collected from more open-ended tasks. Although correctness and diversity are less clearly defined for code generation, SSFT instills diverse and specialized reasoning modes, and RL can quickly leverage these to find optimal forking path for each coding task. In fact, the $\mathrm{Pass@}1$ gain on LCB after SSFT with code traces is slightly larger than the gain on math tasks after SSFT with math traces (Table 1). One explanation is that open-ended tasks involve a much larger search space, and global forking tokens help the model avoid getting stuck in suboptimal reasoning modes.
Table 5: Comparison of SSFT and baselines on one coding benchmark, and three math reasoning tasks that are out-of-distribution relative to the fine-tuning set, since the fine-tuning dataset code1k-4mixed-reasoning contains only coding questions and traces. All methods use Qwen2.5-32B-Instruct as the base model. SFT-mixed-distill denotes standard SFT where different traces for the same prompt are treated as independent training examples. SSFT (random $\bm{\sigma}$) denotes SSFT with a random bipartite matching at each training step. All the three models have gone additional small-scale GFPO for optimal forking selection with 1000 coding questions from Intellect-2-RL. At test time, <think i>  is sampled by the models.
LCB(v5) $\text{AIME 2024}^{*}$ $\text{AIME 2025}^{*}$ $\text{MATH-500}^{*}$
Pass@1: avg over n=16 for LCB, n=32 for math tasks. (* denotes out-of-distribution relative to fine-tuning)
Qwen2.5-32B-Instruct $\triangle$ 23.35 15.80 10.40 80.40
SFT-mixed-distill-32B-code $\bigstar$ 47.13 34.69 24.17 89.39
SSFT-32B-code (random $\bm{\sigma}$) $\bigstar$ 45.36 39.06 31.56 89.46
SSFT-32B-code $\bigstar$ 52.07 43.23 32.82 89.96
Pass@1 of Cons@6: Average performance of majority voting with 6 parallel generations
SFT-mixed-distill-32B-code $\bigstar$ - 41.52 30.30 91.70
SSFT-32B-code (random $\bm{\sigma}$) $\bigstar$ - 46.97 38.18 92.43
SSFT-32B-code $\bigstar$ - 51.82 37.88 93.25
SFT-mixed-distill-32B-code $\bigstar$ - 46.67 33.33 91.40
SSFT-32B-code (random $\bm{\sigma}$) $\bigstar$ - 46.67 43.33 92.00
SSFT-32B-code $\bigstar$ - 56.67 46.67 93.00
A.5 Robustness to Model Family and Size
Setup. We use Qwen3-4B-Base and Llama3.1-8B-Instruct as additional base models, repeating the main experiment from Section 3. Results are reported in Tables 6 and 7, respectively.
Results. Table 6 shows that SSFT remains robust on a smaller 4B model from the newer Qwen3 family (Yang et al., 2025a), and Table 7 shows that it is also robust on an 8B model from the older Llama3.1-8B-Instruct family, achieving higher scores on most reported benchmarks. However, the performance gains are smaller than those observed with the larger Qwen2.5-32B.
Table 6: Comparison of SSFT with baselines on three math reasoning tasks, where all methods use Qwen3-4B-Base as the base model and are trained on the s1k-4mixed-reasoning. SFT-mixed-distill denotes standard SFT where different traces for the same prompt are treated as independent training examples. SSFT (random $\bm{\sigma}$) denotes SSFT with a random bipartite matching. For $\mathrm{Pass@}1$, SSFT uses a graph-based heuristic: it prompts the forking token, <think i> , with the largest number of connected edges in the learned matchings.
AIME 2024 AIME 2025 MATH-500 $\text{LCB(v5)}^{*}$
Pass@1: Average performance of individual generations. LCB is OOD relative to the fine-tuning data.
[…]
Cons@6: Average performance of majority voting with 6 parallel generations
[…]
Table 7: Comparison of SSFT with baselines on three math reasoning tasks, where all methods use Llama3.1-8B-Instruct as the base model and are trained on the s1k-4mixed-reasoning. SFT-mixed-distill denotes standard SFT where different traces for the same prompt are treated as independent training examples. SSFT (random $\bm{\sigma}$) denotes SSFT with a random bipartite matching. For $\mathrm{Pass@}1$, SSFT uses a graph-based heuristic: it prompts the forking token, <think i> , with the largest number of connected edges in the learned matchings; as a result, <think 4> is selected for SSFT-Llama3.1-8B and <think 1> for SSFT-Llama3.1-8B (random $\bm{\sigma}$).
AIME 2024 AIME 2025 MATH-500
Model Pass@1 Cons@32 Pass@1 Cons@32 Pass@1 Cons@32
Llama3.1-8B-Instruct 3.65 13.33 0.94 6.67 47.53 65.40
SFT-mixed-distill-Llama3.1-8B 4.90 13.33 4.90 10.00 61.95 80.00
SSFT-Llama3.1-8B (random $\bm{\sigma}$) 6.77 13.33 4.79 10.00 62.22 79.60
SSFT-Llama3.1-8B 6.98 16.66 5.79 20.00 62.04 80.00
A.6 Generalization to Domains Unseen During SSFT.
$\text{SSFT using math traces}\rightarrow\text{Coding}.$ We follow Wang et al. (2025b) and study out-of-distribution generalization to code generation using LiveCodeBench(V5) as the evaluation task after fine-tuning with only math data s1k-4mixed-reasoning, 1.28k questions from DAPO-Math-17k. In Tables 1 and 4, we see that SSFT improves OOD generalization by about $5\%$ over the baselines and about $19\%$ over the base model in the 32B setting. For the 4B model, Table 6 shows that it is about $1\%$ lower than one baseline, but still improves over the base model by about $7\%$.
$\text{SSFT using coding traces}\rightarrow\text{Math}.$ Since we fine-tuned models with only code generation traces, we also study the other direction. AIME24/25 and MATH-500 are used as OOD tasks. Table 5 shows that SSFT yields significant improvements over the baselines across all math tasks, indicating that the advantage on LCB transfers to math reasoning. 5
A.7 Extended Related Work.
Aggarwal and Welleck (2025); Zhang et al. (2025) are seminal works that use RL to learn a policy controlling how long a reasoning model thinks. During RLVR, Aggarwal and Welleck (2025) appends an explicit “Think for $n$ tokens” instruction to the input prompt and adds a length penalty based on the exact deviation. Zhang et al. (2025) instead appends a higher-level length description and adjusts the length penalty accordingly. We list the differences from our method below, though we believe our method is complementary to many of these RL works, since RLVR requires a strong base model (Yue et al., 2025), which SSFT aims to provide.
Our focus on diversity and correctness vs. their focus on efficiency.  They first enforce a relatively small token budget, $\leq 4096$, and then study accuracy. Our work instead aims to address the plateaued or lower performance caused by overthinking. Even if one particular reasoning mode does not improve with longer CoT, other reasoning modes can provide distinct and still potentially long CoT. We use the distribution of thinking counts as an unbiased metric, in addition to accuracy differences, to assess diversity. We think length is a good diversity indicator for us because SSFT does not explicitly impose any length regularization. Controlling reasoning length through global forking tokens is an emergent behavior, in contrast to these RL works that use explicit length penalties in their objectives.
Experiment settings. Our SFT-OpenR1-93k-7B is a strong enough baseline scoring $46.15\%$ on AIME24, compared to $26\%$ by L1-Exact-7B (Aggarwal and Welleck, 2025), and $30\%$ by TLDR-7B-4k (Aggarwal and Welleck, 2025). Part of the difference likely comes from setup asymmetries: our baseline does not constrain the token limit and uses teacher distillations, whereas these works rely on DeepSeek-R1-Distill-7B, a base model that has gone through SFT using private data. So we would need to run a non-trivial number of experiments to first verify whether their length-controlled optimization remains competitive at much larger token budgets (at least 15,000), and whether their results hold with a base SFT model fine-tuned with public data. Even then, the research question would be whether their RL methods can make SSFT unnecessary, so we would need to compare SSFT plus their method against SFT plus their method. We plan to investigate the synergy between an SSFT base model and different RL methods in future work.
A.8 Algorithm: SSFT Implementation
Algorithm 1 presents the core SSFT implementation with optimal bipartite matching. In practice, the nested-loop computation used to populate $\mathrm{C}$ is fully vectorized and can be executed in a single forward pass. This does not blow up VRAM because (i) we do not store activations for these cost evaluations (no backprop through matching costs), and (ii) We only need to use the first ${\mathrm{T}_{\mathrm{L}}}<{\mathrm{T}_{{\mathbf{r}}}}$ NTP losses to compute the matching cost, as the NTP loss over the first few thousand tokens can already differentiate many reasoning traces in terms of their modes. Nevertheless, our code also supports matching over the full $\mathrm{T}_{{\mathbf{r}}}$} tokens by chunking the computation into a few batches, so this step does not become a VRAM bottleneck. Fine-tuning on 1k questions with 4 traces each, SSFT (optimal matching) took 6.5 h for 6 epochs, compared to 6.1 h for standard SFT, adding only a small overhead.
To support training with variable-size parallel generations, we propose a simple and scalable implementation that incurs no additional VRAM overhead. Rather than concatenating diverse reasoning traces along the sequence dimension, which would require complex sequence parallelism for memory efficiency, our algorithm expands variable-size parallel generations along the batch dimension under distributed training (Appendix A.8).
The primary VRAM bottleneck in SSFT remains the backpropagation Step 8, regardless of whether we use optimal or random matching, because the effective batch size scales with $\mathrm{M}$. To mitigate this, we split the backward pass into several gradient-accumulation steps. Although our experiments use the same number of reasoning traces per question, we also support variable number of targets using our queue-based batching in Algorithm 2. The complication arises when using distributed training with a variable-sized batch, as different processes require the same per-device batch size to perform collective operations. This is mitigated by padding with PAD” sequences to align batch sizes. Our implementation minimizes the number of “PAD” sequences by storing a variable number of targets in a queue and dequeuing multiple items to form a per-device global batch, so smaller batches can be stitched together instead of always being padded.